# RhythmFormer: Extracting rPPG Signals Based on Hierarchical Temporal Periodic Transformer

### v1

@@ Abstract
Remote photoplethysmography (rPPG) is a non-contact method for detecting physiological signals based on facial videos, holding high potential in various applications such as healthcare, affective computing, anti-spoofing, etc. Due to the periodicity nature of rPPG, the long-range dependency capturing capacity of the Transformer was assumed to be advantageous for such signals. However, existing approaches have not conclusively demonstrated the superior performance of Transformer over traditional convolutional neural network methods, this gap may stem from a lack of thorough exploration of rPPG periodicity. In this paper, we propose RhythmFormer, a fully end-to-end transformer-based method for extracting rPPG signals by explicitly leveraging the quasi-periodic nature of rPPG. The core module, Hierarchical Temporal Periodic Transformer, hierarchically extracts periodic features from multiple temporal scales. It utilizes dynamic sparse attention based on periodicity in the temporal domain, allowing for fine-grained modeling of rPPG features. Furthermore, a fusion stem is proposed to guide self-attention to rPPG features effectively, and it can be easily transferred to existing methods to enhance their performance significantly. RhythmFormer achieves state-of-the-art performance with fewer parameters and reduced computational complexity in comprehensive experiments compared to previous approaches. The codes are available at https://github.com/zizheng-guo/RhythmFormer.
@@ 1 Introduction
Figure 1: Sparsity in attention due to periodicity. Given the red block as a Query, the global spatio-temporal correlations are computed. Brighter regions in the attention score and key responses in the frame indicate stronger attention. From the figure, it is evident that the brightness of the key responses in the frame is predominantly concentrated in the skin area, while the brightness in the attention score is mainly focused on the troughs (regions with the same phase as the Query). The attention scores exhibit a notable sparsity.
@@
Blood Volume Pulse (BVP) is an important physiological signal, further allowing for the extraction of key vital signs such as heart rate (HR) and heart rate variability (HRV). Photoplethysmography (PPG) is a non-invasive monitoring method that utilizes optical means to measure changes in blood volume within living tissues. The physiological mechanism underlying BVP extraction from PPG is rooted in variations in light absorption and scattering due to changes in blood volume during cardiac contraction and relaxation in subcutaneous blood vessels. These variations result in periodic color signal alterations on imaging sensors, imperceptible to the human eye [45, 3]. Traditionally, BVP extraction required the use of contact sensors, introducing inconveniences and limitations. In recent years, non-contact methods for extracting BVP information, specifically remote photoplethysmography (rPPG), have garnered increasing attention [29, 12, 28].
@@
Early research on rPPG primarily relied on signal processing based methods to analyze skin, typically facial, videos for extracting periodic signals [45, 34, 19, 9, 10, 33, 21, 47, 48]. However, due to the weak amplitude of rPPG signals and the presence of strong noise from environmental factors like lighting and motion, achieving high accuracy solely through signal processing in complex environments is challenging. With the advancement of deep learning, there is a growing trend towards data-driven methods. In recent years, significant attention has been directed towards the transformer within the rPPG domain. Compared to CNNs, transformers excel in capturing long-range dependency. This is particularly advantageous for extracting BVP signals characterized by quasi-periodicity and simultaneous existence across all skin regions. However, existing research has not convincingly demonstrated that transformers outperform traditional CNNs. This may be attributed to the quadratic scaling exhibited by Transformers with sequence length [1], a challenge further compounded by the high dimensionality of video. Past work typically transformed inputs into coarse-grained tokens, and due to artifacts and head movement [56, 57], these coarse tokens inevitably encompass both regions rich in rPPG information and regions with significant noise. The introduction of noise through coarse-grained extraction, combined with the inherently weak nature of rPPG signals, results in a low signal-to-noise ratio.
To address the aforementioned issues, we observe that: as shown in Figure 1, spatially, numerous regions (including the background and areas without exposed skin) are irrelevant to rPPG information. Temporally, the attention of rPPG displays notable temporal sparsity. We anticipate attention to concentrate on skin regions with similar temporal phases, making a substantial portion of self-attention computations invaluable. Based on the observed characteristics, we propose RhythmFormer, which explicitly leverages the periodicity of rPPG. RhythmFormer explores periodic features at multiple temporal scales, utilizing the temporal sparsity induced by periodicity. Initially, computing spatio-temporal correlations within a large spatio-temporal receptive field aims to aggregate regions with high phase similarity temporally and rich rPPG signals spatially. Subsequently, finer-grained feature extraction is conducted from these high-correlation regions. Concurrently, in addressing the challenge of pronounced self-attention deviation due to the inherently weak nature of rPPG features, we devised the fusion stem. This design allows frame-level representations to be aware of BVP wave variations, effectively guiding the self-attention mechanism.
@@
1.We propose RhythmFormer, a fully end-to-end method that explicitly leverages periodicity to design the periodic sparse attention mechanism, which finely extracts periodic rPPG features across multiple temporal scales.
2.We propose a plug-and-play fusion stem module, enabling frame-level representation awareness of BVP wave variations. This module can be seamlessly integrated into other methods, thereby significantly improving the performance of existing approaches.
3.We conduct extensive experiments on both intra-dataset and cross-dataset scenarios. Results demonstrate that our method outperforms previous state-of-the-art methods in comprehensive testing with fewer parameters and reduced computational requirements.
Figure 2: Framework of RhythmFormer. It consists of fusion stem, patch embedding, Hierarchical Temporal Periodic Transformer, and rPPG predictor head. The proposed Hierarchical Temporal Periodic Transformer consists of three stages and each stage contains a TPT Block with a sampling coefficient of n, where n is the number of the sample block.
@@ 2.1 Remote Physiological Estimation
Previous research on rPPG predominantly relied on traditional signal processing methods to analyze facial videos for extracting their periodic signals [45, 34, 19, 9, 10, 21, 47, 48, 33]. With the advancement of deep learning, an increasing number of data-driven approaches have emerged, showcasing a trend in backbones transitioning from 2D CNNs [39, 30, 31, 3, 32, 22, 6] to 3D CNNs [51, 52, 42, 60, 18, 20] and further to transformer [54, 15, 36, 56, 57, 23, 13, 59, 37]. The superiority of 3D CNNs over 2D CNNs arises from their enhanced capability to capture temporal information for tasks like rPPG. In comparison to 3D CNNs, transformers excel in capturing long-range spatial and temporal correlations, making them particularly advantageous for extracting quasi-periodic PPG signals. However, existing transformer-based methods are either non-end-to-end, requiring a complex preprocessing procedure and potentially overlooking global contextual cues outside predefined ROIs [13, 59, 37, 15], or they are based on Video Transformer, tending to extract features at a coarse granularity [36, 56, 57], or they are based on Vision Transformer, which may disregard long-range temporal relationships [23]. None of these approaches has demonstrated a significant performance advantage over CNN methods.
2.2 Sparse Attention Mechanisms
@@
The conventional attention mechanism poses a significant computational burden and demands substantial memory resources as it involves pair-wise token interaction across all spatial-temporal locations [44, 1]. Numerous solutions have been proposed to address this issue [5, 46, 8]. Sparse attention has gained recent popularity, particularly fueled by the remarkable success of the Swin Transformer [25]. Subsequent works have introduced various hand-crafted sparse attention, such as expanding windows [43, 49] or cross-shaped windows [11]. Some endeavors aim to adapt sparse patterns to data, exemplified by DAT [50], TCFormer [58], and DPT [4]. Biformer [61] explores a dynamic, query-aware sparse attention mechanism, introducing new possibilities for sparse attention. Previous rPPG-related works [56, 57, 23, 13] primarily employed coarse-grained global attention without incorporating sparse attention mechanisms. This might be attributed to two potential reasons: One is that rPPG information is present in the skin regions across all temporal periods, making it difficult to achieve sparse attention through hand-crafted static attention mechanisms; Another is that most of the sparse attention mechanisms mentioned above are designed for use on two-dimensional images, posing a challenge in their application to three-dimensional video data.
@@ 3 Methodology
In Section 3.1, the general framework of RhythmFormer will be presented, followed by the Fusion Stem in Section 3.2 and the Hierarchical Temporal Periodic Transformer (TPT) in Section 3.3. Lastly, Section 3.4 introduces the proposed HR hybrid loss.
@@ 3.1 The general framework of RhythmFormer
As shown in Figure 2, RhythmFormer primarily consists of a fusion stem, the patch embedding, a Hierarchical Temporal Periodic Transformer, and an rPPG prediction head (MLP). We utilize the fusion stem to integrate the difference frame with the raw frame, enabling the capability to be aware of frame-level rPPG information. This enhances the representation features and guides the transformer to focus on the rPPG signal. Specifically, given an RGB video input $X\in\mathbb{R}^{3\times T\times H\times W}$, $X_{stem}=fusion\_stem(X)$, where $X_{stem}\in\mathbb{R}^{C\times T\times H/4\times W/4}$, and C, T, W, H indicate channel, sequence length, width, and height, respectively.
@@
Subsequently, the output of patch embedding will be fed into the Hierarchical Temporal Periodic Transformer, extracting rPPG features at a fine granularity from different time scales through periodic sparse attention mechanisms. Specifically, this involves three TPT Blocks, where the input and output of the $i_{th}$ block are denoted as $X_{block\_i}\in R^{C\times T\times H/16\times W/16}$. Finally, the rPPG features will be projected into BVP waves through a prediction head.
@@ 3.2 Fusion Stem
Handling the conflict between strong noise and weak features is a key issue in the rPPG task. Some non-end-to-end methods [27, 32, 12, 13, 37] use color space transformations based on ROI regions to obtain spatio-temporal signal maps. However, these methods ignore global contextual cues outside the predefined ROIs and require strict preprocessing steps, making them less suitable for deployment on mobile devices. In previous end-to-end rPPG research, there are primarily two ways to take video input. The first one is to directly use raw frames as input [51, 56, 57]. The second one is to input the normalized frame difference, while also inputting the raw frames to provide attention. The difference frame is guided by the original raw frame to focus on the skin region [3, 22, 60]. The difference frame holds an advantage as it effectively removes the influence of fixed component noise (e.g., skin color and camera noise). However, other noise sources such as ambient light and motion artifacts still exist. This low signal-to-noise ratio poses a significant challenge for transformer models based on self-attention mechanisms. When using raw frames as input, self-attention struggles to focus on the subtle changes in skin color compared to more prominent facial features. There is a higher correlation between similar spatial regions in each frame, rather than between temporally phase-similar regions. Consequently, a significant proportion of the self-attention scores originate from facially similar regions rather than capturing the rPPG features manifested by skin color changes, leading to a substantial deviation between the actual focus and the desired ones.
@@
Based on this, we propose the fusion stem that integrates the difference frame into the raw frame, enabling frame-level representation to be aware of BVP wave variations. This fusion stem effectively enhances rPPG features with minimal additional computational cost. The inclusion of this module not only guides the self-attention mechanism of the transformer to reinforce its focus on rPPG representation but also achieves outstanding performance across all end-to-end methods [see Section 4.5].
@@
As shown in Figure 2, for an input video $X\in\mathbb{R}^{3\times T\times H\times W}$, first temporal shift is applied to obtain $X_{t-2}$, $X_{t-1}$, $X_{t}$, $X_{t+1}$ and $X_{t+2}$. Then, the differences between consecutive frames are computed in chronological order, resulting in $D_{-2}$, $D_{-1}$, $D_{1}$ and $D_{2}$. These difference frames, along with the raw frames, are passed through Stem_1 for primary feature extraction. Stem_1 consists of a 2D convolution with a kernel size of $(5\times 5)$, cascaded by batch normalization(BN), ReLU, and MaxPool.
$\small X_{diff}\ =Stem\_1(Concat\left(D_{-2},D_{-1},D_{1},D_{2}\right))$ (1)
$\small X_{origin}\ =Stem\_1(X_{t})$ (2)
@@
Then the features of the difference video and the raw video are merged and the feature representation is further enhanced through Stem_2, which also consists of a 2D convolution with a kernel size of $(3\times 3)$, cascaded by BN and ReLU.
$\small X_{stem}$ $=\alpha\ \ast\ Stem\_2(X_{origin})+$ (3)
$\beta\ \ast\ Stem\_2(\alpha\ \ast\ X_{origin}+\beta\ \ast\ X_{diff})$
The fusion coefficients $\alpha$ and $\beta$ are both set to 0.5 experimentally.

### v3

# RhythmFormer: Extracting Patterned rPPG Signals based on Periodic Sparse Attention

@@ Abstract
Remote photoplethysmography (rPPG) is a non-contact method for detecting physiological signals based on facial videos, holding high potential in various applications. Due to the periodicity nature of rPPG signals, the long-range dependency capturing capacity of the transformer was assumed to be advantageous for such signals. However, existing methods have not conclusively demonstrated the superior performance of transformers over traditional convolutional neural networks. This may be attributed to the quadratic scaling exhibited by transformer with sequence length, resulting in coarse-grained feature extraction, which in turn affects robustness and generalization. To address that, this paper proposes a periodic sparse attention mechanism based on temporal attention sparsity induced by periodicity. A pre-attention stage is introduced before the conventional attention mechanism. This stage learns periodic patterns to filter out a large number of irrelevant attention computations, thus enabling fine-grained feature extraction. Moreover, to address the issue of fine-grained features being more susceptible to noise interference, a fusion stem is proposed to effectively guide self-attention towards rPPG features. It can be easily integrated into existing methods to enhance their performance. Extensive experiments show that the proposed method achieves state-of-the-art performance in both intra-dataset and cross-dataset evaluations. The codes are available at https://github.com/zizheng-guo/RhythmFormer.
keywords:
Remote physiological measurement , Periodic sparse attention
journal: arXiv
[…]
@@ 1 Introduction
Blood Volume Pulse (BVP) is an important physiological signal, further allowing for the extraction of key vital signs such as heart rate (HR) and heart rate variability (HRV). Photoplethysmography (PPG) is a non-invasive monitoring method that utilizes optical means to measure changes in blood volume within living tissues. The physiological mechanism underlying BVP extraction from PPG is rooted in variations in light absorption and scattering due to changes in blood volume during cardiac contraction and relaxation in subcutaneous blood vessels. These variations result in periodic color signal alterations on imaging sensors, imperceptible to the human eye [1, 2]. Traditionally, BVP extraction required the use of contact sensors [3, 4], introducing inconveniences and limitations. In recent years, non-contact methods for extracting BVP information, specifically remote photoplethysmography (rPPG), have garnered increasing attention [5, 6]. These methods possess significant research value not only within the field of healthcare but also hold broad application prospects in a multitude of fields, including affective computing, deepfake detection, and anti-spoofing, among others [7].
Figure 1: Sparsity in attention due to periodicity. Given the red block as a Query, the global spatio-temporal correlations are computed. Brighter regions in the attention score (top row) and key response (bottom row) in the frame indicate stronger attention. From the figure, it is evident that the brightness of the key response in the frame is predominantly concentrated in the facial area, while the brightness in the attention score is mainly focused on the troughs (regions with the same phase as the Query). The attention scores exhibit a notable sparsity.
@@
Early research on rPPG primarily relied on signal processing-based methods to analyze facial videos for extracting periodic signals [1, 8, 9]. However, due to the weak nature of rPPG signals, which have low amplitude and are subject to strong noise and artifacts from environmental factors such as lighting and motion, achieving high accuracy solely through signal processing in complex environments is challenging. With the advancement of deep learning, there is a growing trend towards data-driven methods. In recent years, significant attention has been directed towards the transformer within the rPPG field. Compared with convolutional neural networks (CNNs), transformers excel in capturing long-range dependency. This enables them to model BVP features by leveraging intra-frame (facial region) and inter-frame (similar temporal phase) correlations, making them highly advantageous for quasi-periodic rPPG signals. However, existing research has not convincingly demonstrated that transformers outperform traditional CNNs. This may stem from their merely passive exploitation of the long-range dependency advantages for periodicity, which is achieved through global attention or static sparse attention. Nevertheless, the quadratic scaling exhibited by transformers with sequence length [10], further compounded by the high dimensionality of video, restricted these approaches to transforming inputs into coarse-grained tokens. Due to artifacts and head movement, these coarse tokens inevitably encompass both regions rich in rPPG information and regions with significant noise [11]. The introduction of noise through coarse-grained extraction, combined with the inherent weak nature of rPPG signals, results in a low signal-to-noise ratio, significantly impacting robustness and generalization.
To address this challenge, we focus on the phenomenon of notable sparsity in temporal attention due to periodicity, as shown in Figure 1. Inspired by findings in cognitive psychology related to pre-attentive vision, which refers to the processing of visual stimuli before they are selected by attention for more complete analysis [12, 13], we propose a periodic sparse attention mechanism. This mechanism explicitly leverages the core periodic nature of rPPG, integrating it into the attention design. As illustrated in Figure 2, a pre-attention stage is introduced before the conventional attention mechanism. The pre-attention stage learns potential periodic patterns in the rPPG signals and involves temporally aggregating regions with high phase similarity and spatially aggregating areas rich in rPPG signals, effectively filtering out numerous irrelevant attention computations. This allows subsequent attention to focus on fine-grained feature extraction.
Moreover, the inherent weak nature of rPPG signals and the susceptibility of fine-grained features to noise interference collectively lead to noticeable self-attention biases. To address this challenge, we devise a fusion stem. This design allows frame-level representations to capture BVP wave variations, effectively guiding the self-attention mechanism.
Figure 2: The proposed periodic sparse attention mechanism utilizes the periodic patterns learned through pre-attention to gather high-score regions (facial regions with similar temporal phases), which filters out a substantial number of irrelevant attention computations. This allows for finer-grained token extraction across dimensions H, W, and T compared with vanilla attention.
@@
$\bullet$ Leveraging the sparsity of temporal attention stemming from periodicity, we design a periodic sparse attention mechanism, which filters out redundant computations through pre-attention, enabling fine-grained attention.
$\bullet$ We propose a plug-and-play fusion stem module, enabling frame-level representation awareness of BVP wave variations. This module can be seamlessly integrated into other methods, significantly improving the performance of existing approaches.
$\bullet$ We propose RhythmFormer. Extensive experiments conducted on both intra-dataset and cross-dataset scenarios demonstrate that the proposed method outperforms previous state-of-the-art approaches.
@@ 2.1 Remote Physiological Estimation
Early research on rPPG predominantly relied on traditional signal processing methods to analyze facial videos for extracting quasi-periodic signals [1, 14, 8, 9]. With the advancement of deep learning, an increasing number of data-driven approaches have emerged, showcasing a trend in encoder transitioning from 2D CNNs [15, 16, 2, 17, 18] to 3D CNNs [19, 20, 21, 22] and further to transformer [11, 23, 24]. For video-based time-series tasks such as rPPG, 3D CNNs offer a significant advantage over 2D CNNs in the extraction of temporal information. Compared to 3D CNNs, transformers excel in capturing long-range spatial and temporal correlations, which enables them to model BVP signals by leveraging intra-frame (skin region) and inter-frame (similar temporal phase) correlations. However, many of the most recent self-supervised methods or applications still employ CNN architectures as encoders [21, 22, 18]. For example, Greip [18] uses ResNet as the encoder through the integration of explicit and implicit prior knowledge to improve the generalization performance of the model. This may stem from the fact that transformers have not shown a significant performance advantage over CNNs, while CNNs are more user-friendly. The theoretical advantages of transformers have not been fully realized, potentially due to the constraints imposed by quadratic complexity in previous transformer-based approaches. These constraints have led to a coarse-grained extraction of features, thereby not fully leveraging the transformers’ capacity for capturing dependencies.
Faced with the challenge of quadratic complexity, the proposed RhythmFormer not only passively leverages periodic advantages for long-range inter-frame attention, as in previous work, but also proactively exploits periodicity at a deeper level. It introduces a pre-attention stage that filters out a substantial amount of irrelevant attention computations through the learning of periodic patterns, thereby achieving more fine-grained attention at a lower computational cost.
2.2 Sparse Attention Mechanism
@@
The conventional attention mechanism poses a significant computational burden and demands substantial memory resources as it involves pair-wise token interaction across all spatial-temporal locations [25, 10]. As an effective solution, sparse attention has gained recent popularity [26, 27, 28]. Previous rPPG works [11, 23, 24] primarily employed coarse-grained global attention without incorporating sparse attention mechanisms. This might be attributed that rPPG information is present in the skin regions across all temporal periods, making it difficult to achieve sparse attention through hand-crafted static attention mechanisms.
In contrast to global attention or static sparse attention in previous work, we embed the quasi-periodic nature of rPPG into the sparse attention mechanism. Based on the sparsity of temporal attention stemming from periodicity, we design a periodic sparse attention mechanism tailored to periodic tasks. This mechanism leverages periodicity to filter out a significant amount of redundant computations, enabling fine-grained sparse attention.
Figure 3: The general framework of RhythmFormer. It consists of fusion stem, patch embedding, hierarchical temporal periodic transformer, and rPPG predictor head. The hierarchical temporal periodic transformer consists of three temporal periodic transformer (TPT) Blocks and two temporal downsampling modules between TPT Blocks. The attention mechanism in the TPT block is periodic sparse attention. It introduces a pre-attention stage before the vanilla attention, which selects the top-k tokens to filter out irrelevant key (K) and value (V).
@@ 3 Methodology
In Section 3.1, the general framework of RhythmFormer is presented, followed by the fusion stem in Section 3.2 and the hierarchical temporal periodic transformer in Section 3.3. Finally, the model training is introduced in Section 3.4.
@@ 3.1 The General Framework of RhythmFormer
As shown in Figure 3, RhythmFormer consists of fusion stem, patch embedding, hierarchical temporal periodic transformer, and rPPG prediction head (Multilayer Perceptron, MLP). We utilize the fusion stem to integrate the difference frame with the raw frame, enabling the capability to capture frame-level rPPG information. This enhances the feature representation and guides the transformer to focus on the rPPG signal. Specifically, given an RGB video input $X\in\mathbb{R}^{3\times T\times H\times W}$, $X_{stem}=fusion\_stem(X)$, where $X_{stem}\in\mathbb{R}^{C\times T\times H/4\times W/4}$, and C, T, W, H indicate channel, sequence length, width, and height, respectively.
@@
Subsequently, the output of patch embedding will be fed into the hierarchical temporal periodic transformer, extracting rPPG features at a fine granularity from different temporal scales through periodic sparse attention modules. It consists of three TPT Blocks with a downsampling factor of n. Depending on the different n, there is the temporal downsampling module before the TPT block. The input and output of the hierarchical temporal periodic transformer are denoted as $X_{tptin},X_{tptout}\in R^{C\times T\times H/16\times W/16}$. Finally, the rPPG features will be projected into BVP waves through the predictor head.
@@ 3.2 Fusion Stem
Handling the conflict between strong noise and weak features is a key issue in the rPPG task. Some methods [16, 30, 31, 18] use color space transformations based on ROI regions to obtain spatiotemporal signal maps. However, these methods ignore global contextual cues outside the predefined ROIs and require strict preprocessing steps. There are also some methods that take raw video as input, mainly categorized into two ways. The first one is to directly use raw videos as input [19, 11, 32]. The second one is to input the normalized frame differences, while also inputting the raw frames to provide attention. The frame differences are guided by the raw frames to focus on the skin region [2, 17, 20]. The frame differences hold an advantage as it effectively removes the influence of fixed component noise (e.g., skin color and camera noise). However, other artifacts such as ambient light variation and motion artifacts still exist. This low signal-to-noise ratio poses a significant challenge for transformers based on self-attention mechanisms. When using raw frames as input, self-attention struggles to focus on the subtle changes in skin color compared with more prominent facial features. There is a higher correlation between similar spatial regions in each frame, rather than between temporally phase-similar regions. Consequently, a significant proportion of the self-attention scores originate from facially similar regions rather than capturing the rPPG features manifested by skin color changes, leading to a substantial deviation between the actual focus and the desired ones.
@@
Based on this, we propose the fusion stem that integrates the frame differences into the raw frames, enabling frame-level representation to capture BVP wave variations. The fusion stem effectively enhances rPPG features at an extremely low computational cost, guiding the self-attention mechanism to reinforce the focus on rPPG representation. It can also significantly improve the performance of existing methods in a plug-and-play manner (see ablation study in Section 4.6).
The fusion stem is inspired by the short-term TDM [33], but differ in motivation and implementation details. Firstly, TDM is a low-resolution architecture designed to fuse prominent motion patterns. Conversely, our fusion stem employs a fine-grained extraction approach to capture subtle inter-frame rPPG signals. Secondly, TDM utilizes residual blocks to extract motion features and integrate them into the raw frames, whereas fusion stem primarily focuses on fusing inter-frame differences without adding dedicated feature extraction modules. Instead, it employs the existing convolutions in the vanilla stem to downsample and perform preliminary feature extraction.
@@
Specifically, as shown in Figure 3, for an input video $X\in\mathbb{R}^{3\times T\times H\times W}$, the temporal shift is applied to obtain $X_{t-2}$, $X_{t-1}$, $X_{t}$, $X_{t+1}$ and $X_{t+2}$. Then, the differences between consecutive frames are computed in chronological order, resulting in $D_{t-2}$, $D_{t-1}$, $D_{t+1}$ and $D_{t+2}$. These frame differences, along with the raw frames, are passed through two separate instances of ${Stem}_{1}$ for primary feature extraction. ${Stem}_{1}$ consists of a 2D convolution layer with a kernel size of $(5\times 5)$, cascaded by batch normalization (BN), ReLU, and MaxPool. The input dimension is 3 when taking the raw frames as the input and 12 when taking the concatenation of frame differences as the input.
$X_{diff}=Stem_{1}(Concat(D_{t-2},D_{t-1},D_{t+1},D_{t+2})),$ (1)
$X_{raw}=Stem_{1}(X_{t}),$
$X_{fusion}=X_{raw}+X_{diff}.$
@@
Then, the frame difference feature and the fused feature representation are further enhanced through $Stem_{2}$, which also consists of a 2D convolution layer with a kernel size of $(3\times 3)$, cascaded by BN and ReLU.
$X_{stem}=Stem_{2}(X_{fusion})+Stem_{2}(X_{diff}).$ (2)
This design integrates multi-level features extracted from the frame differences of adjacent frames into the raw frame. Compared with the raw frame input, it significantly enhances the rPPG features, guiding the self-attention mechanism. In Figure 4, the visual results of $X_{diff}$ and frame differences are compared in (b) and (c). It can be observed that the frame difference eliminates fixed component noise and focuses on subtle local color changes, but it also leads to the amplification of even the smallest noise signals, resulting in a very low signal-to-noise ratio. In contrast, the fusion stem focuses entirely on extracting frame differences from the face region. This method, which extracts deep features through convolution across adjacent five frames, contains more rPPG information and less noise compared to simple frame differences, greatly mitigating its weaknesses.
(a) $X_{raw}$
(b) $X_{diff}$
(c) Frame Difference
Figure 4: Visualization comparison of features extracted by the fusion stem and frame differences.